OpenStudy (anonymous):

Solve the proportion: 4x/3=15/7 A.28/45 B.45/28 C.18/11 D. 17

OpenStudy (anonymous):

cross multiply and divide

OpenStudy (uri):

4x/3=15/7 x =15 x3/7 x 4 x=45/28 You will do cross multiply and divide.